How long will it take $47,000 placed in a savings account at 10% interest to grow into $72,000?\
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

72,000=47,000*(1.1)^n

(72,000/47,000)=(1.1)^n

Taking log on both sides;

log (72,000/47,000)=n*log 1.1

n=log (72,000/47,000)/log 1.1

=4.48 years(Approx).